Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $0.75 coming 3 years from today. The dividend should grow rapidly - at a rate of 50% per year - during Years 4 and 5, but after Year 5, growth should be a constant 8% per year. If the required return on Computech is 15%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Current stock price is $ 14.92

As per dividend discount model, current stock price is the present value of dividends which is calculated as follows:
Step-1:Present Value of non-constant growth stage dividend
Year Dividend Discount factor Present Value
a b c=1.15^-a d=b*c
3 0.75 0.658 0.49
4 1.13 0.572 0.64
5 1.69 0.497 0.84
Total 1.98
Working:
Dividend of year:
3 = 0.75
4 = 0.75 * 1.50 = 1.13
5 = 1.125 1.50 = 1.69
Step-2:Present value of constant growth stage dividend
Present Value = D5*(1+g)/(Ke-g)*DF5
= 12.94
Where,
D5 = Year 5 dividend = 1.69
g = Growth Rate = 8%
Ke = Required Return = 15%
DF5 = Discount factor of Year 5 = 0.497
Step-3:Calculation of current price of stock
Current Price = Present Value of dividends
= 1.98 + 12.94
= 14.92